# Improve the Precision of Area Under the Curve Estimation for Recurrent Events Through Covariate Adjustment

### v2

@@ Abstract
The area under the curve (AUC) of the mean cumulative function (MCF) has recently been introduced as a novel estimand for evaluating treatment effects in recurrent event settings, capturing a totality of evidence in relation to disease progression. While the Lin-Wei-Yang-Ying (LWYY) model is commonly used for analyzing recurrent events, it relies on the proportional rate assumption between treatment arms, which might be violated in practice. In contrast, the AUC under MCFs does not depend on such proportionality assumptions and offers a clinically interpretable measure of treatment effect. To improve the precision of the AUC estimation while preserving its unconditional interpretability, we propose a nonparametric covariate adjustment approach. This approach guarantees efficiency gain compared to unadjusted analysis, as demonstrated by theoretical asymptotic distributions, and is universally applicable to various randomization schemes, including both simple and covariate-adaptive designs. Extensive simulations across different scenarios further support its advantage in increasing statistical power. Our findings highlight the importance of covariate adjustment for the analysis of AUC in recurrent event settings, offering practical guidance for its application in randomized clinical trials.
@@ 1 Introduction
In recent years, the recurrent event analysis has gained traction in clinical studies, particularly in chronic conditions such as multiple hospitalizations for heart failure, recurrent relapses in multiple sclerosis, or frequent exacerbation in asthma [Solomon et al., 2024, Rogers et al., 2014, Cohen and Rudick, 2011, Keene et al., 2007]. A widely adopted approach for analyzing recurrent event data in clinical trials is the Lin-Wei-Yang-Ying (LWYY) model [Lin et al., 2000]. The LWYY model assumes that the mean cumulative functions (MCFs) between two treatment arms are proportional, allowing treatment effects to be assessed through the ratio of MCFs. However, this approach relies on the proportionality assumption, which implies that the rate ratio is constant over time, similar to the proportional hazards assumption in the Cox model. In practice, this assumption may be violated, particularly in complex scenarios where treatment effects are time-varying. For example, the recent FINEARTS-HF study demonstrated time-dependent variation in rate ratio estimates [Vaduganathan et al., 2024]. Such violations can compromise the clinical interpretability of the LWYY model. Therefore, more flexible estimands that do not rely on such restrictive assumptions are required.
The recent ICH E9(R1) addendum also emphasizes the importance of robust definition quantifying the treatment effects in the settings of recurrent events with minimal assumptions. In this context, the area under the curve (AUC)-based approach for recurrent events proposed by Claggett et al. [2022] has gained considerable attention, offering a clinically meaningful summary of the cumulative burden of recurrent events without relying on the proportionality assumption [Gregson et al., 2023]. By calculating the area under the MCF for each treatment arm, one can derive a summary measure that reflects the overall mean burden of events experienced by patients over the follow-up period. A larger AUC signifies a greater cumulative burden and treatment comparisons based on the difference or ratio of AUCs allow for intuitive interpretation in terms of absolute or relative reductions in disease burden. AUCs can be consistently estimated by Ghosh-Lin estimator irrespective of whether the proportionality assumption holds, making it a robust and interpretable estimand for analyzing recurrent event data [Ghosh and Lin, 2000].
@@ 2.1 Data and estimand
Let $D$ represent the time to a terminal event and write $N^{D}(t)=I(D\leq t)$ as the counting process for $D$. Let $N^{*}(t)$ be the number of recurrent events up to time $t$. Let $C$ denote the independent censoring time. An observation consists of $\{N(\cdot),T,\delta\}$, where $N(t)=N^{*}(t\wedge C)$, $T=C\wedge D$, $\delta=\mathbb{I}(D\leq C)$, and $a\wedge b=\hbox{min}(a,b)$. The observed data set $\mathcal{D}=\{N_{i}(\cdot),T_{i},\delta_{i}\}$, for $i=1,\dots,n$, consists of independent realizations of $\{N(\cdot),T,\delta\}$.
Let $S(u)=\mathbb{P}(D\geq u)$ be the survival function, and $dR(u)=\mathbb{E}\{dN^{*}(u)\mid T\geq u\}=\mathbb{E}\{dN^{*}(u)\mid D\geq u\}$. Then the mean cumulative function (MCF) is defined as
@@
Let $\hat{\mu}(t)$ and $\hat{U}(\tau)$ denote estimates of $\mu(t)$ and $U(\tau)$, obtained by substituting $S(u)$ with the Kaplan-Meier estimator $\hat{S}(u)$ and $dR(u)$ with the Nelson-Aalen estimator $d\hat{R}(u)=\sum_{i=1}^{n}dN_{i}(u)/\sum_{i=1}^{n}\mathbb{I}(T_{i}\geq u)$.
@@
The estimates $\hat{\psi}_{i}(\tau)$, $\hat{P}_{i}$, and $\hat{Q}_{i}$ are obtained by replacing $S(u)$ with $\hat{S}(u)$, $\mu(s)$ with $\hat{\mu}(s)$, and estimating $\mathbb{P}(T\geq u)$ as $\sum_{i=1}^{n}\mathbb{I}(T_{i}\geq u)/n$. The martingales are estimated as: $d\hat{M}_{i}(u)=dN_{i}(u)-\mathbb{I}_{i}(T\geq u)d\hat{R}(u)$, and $d\hat{M}_{i}^{D}(u)=dN_{i}^{D}(u)-\mathbb{I}_{i}(T\geq u)d\hat{A}^{D}(u)$, where $d\hat{A}^{D}(u)=\sum_{i=1}^{n}dN^{D}_{i}(u)/\sum_{i=1}^{n}\mathbb{I}(T_{i}\geq u)$.
Denote $U_{0}(\tau)$ and $U_{1}(\tau)$ as the area under the MCF for placebo and treatment arms. The treatment effect can be quantified as either the difference, $\Delta U(\tau)=U_{1}(\tau)-U_{0}(\tau)$, or the ratio, $U_{1}(\tau)/U_{0}(\tau)$. For convenience in later derivations, we define $R(\tau)=\hbox{log}\frac{U_{1}(\tau)}{U_{0}(\tau)}$, which represents the log-transformed ratio of the areas under the MCF.
2.2 Nonparametric Covariate Adjustment
We extend the nonparametric covariate adjustment approach introduced by Ye et al. [2022, 2023, 2024] to the AUC-based summary. The key concept involves linearizing the test statistic to derive an outcome for each patient, followed by the application of generalized regression adjustment or augmentation to these derived outcomes [Zhang et al., 2008, Tsiatis et al., 2008].
2.2.1 Difference of AUCs
@@
Let $\Delta\hat{U}(\tau)=\hat{U}_{1}(\tau)-\hat{U}_{0}(\tau)$ and $\hat{\mathcal{U}}_{L,\Delta}=\frac{\sqrt{n_{0}}\sqrt{n_{1}}}{n}\left\{\Delta\hat{U}(\tau)-\Delta U(\tau)\right\}$, where $n_{0}$ and $n_{1}$ represent the sample sizes for the placebo and treatment arms, respectively. The linearization of $\hat{\mathcal{U}}_{L,\Delta}$ is as follows:
$\hat{\mathcal{U}}_{L,\Delta}=\frac{1}{n}\sum_{i=1}^{n}\left\{I_{i}(\hat{P}^{\Delta}_{i}-\hat{Q}^{\Delta}_{i})-(1-I_{i})(\hat{P}^{\Delta}_{i}-\hat{Q}^{\Delta}_{i})\right\}+o_{p}(1)$
where $I_{i}=1$ if subject $i$ is in the treatment arm and 0 otherwise, $\hat{P}^{\Delta}_{i}=\left(\frac{n_{1}}{n_{0}}\right)^{\frac{1}{2}(-1)^{I_{i}}}\hat{P}_{i}$ and $\hat{Q}^{\Delta}_{i}=\left(\frac{n_{1}}{n_{0}}\right)^{\frac{1}{2}(-1)^{I_{i}}}\hat{Q}_{i}$. The unadjusted test statistic on the difference of AUCs is $\mathcal{T}_{L,\Delta}=\sqrt{n}\hat{\mathcal{U}}_{L,\Delta}/\hat{\sigma}_{L,\Delta}$, where
$\hat{\sigma}_{L,\Delta}^{2}=\frac{1}{n}\left[n_{0}\mathbb{E}\left\{\hat{\psi}_{1}(\tau)^{T}\hat{\psi}_{1}(\tau)\right\}+n_{1}\mathbb{E}\left\{\hat{\psi}_{0}(\tau)^{T}\hat{\psi}_{0}(\tau)\right\}\right]$
Here, $\hat{\psi}_{1}(\tau)$ and $\hat{\psi}_{0}(\tau)$ consist of $\hat{\psi}_{i}(\tau)$ from subjects in the treatment and placebo arms, respectively. Derivations of $\hat{\mathcal{U}}_{L,\Delta}$ and the asymptotic properties of $\mathcal{T}_{L,\Delta}$ are provided in detail in the Section A.1 of the Supplementary Materials.
Let $\boldsymbol{X}^{*}$ be a $p$-dimensional vector of covariates. Define $X_{i}=X_{i}^{*}-\bar{X}$, where $X_{i}^{*}$ is the $i$th column of $\boldsymbol{X}^{*}$, and $\bar{X}$ is the sample mean of all the $X_{i}^{*}$, ensuring $\mathbb{E}(X_{i})=0$. Following Ye et al. [2024], we treat $\hat{P}^{\Delta}_{i}$ and $\hat{Q}^{\Delta}_{i}$ as outcomes and apply the generalized regression adjustment or augmentation [Tsiatis et al., 2008]. This results in the covariate-adjusted $\hat{\mathcal{U}}_{CL,\Delta}$:
$\hat{\mathcal{U}}_{CL,\Delta}=\frac{1}{n}\sum_{i=1}^{n}\Bigg{[}I_{i}\left\{(\hat{P}^{\Delta}_{i}-X_{i}^{T}\hat{\beta}_{1,P}^{\Delta})-(\hat{Q}^{\Delta}_{i}-X_{i}^{T}\hat{\beta}_{1,Q}^{\Delta})\right\}-(1-I_{i})\left\{(\hat{P}^{\Delta}_{i}-X_{i}^{T}\hat{\beta}_{0,P}^{\Delta})-(\hat{Q}^{\Delta}_{i}-X_{i}^{T}\hat{\beta}_{0,Q}^{\Delta})\right\}\Bigg{]}=\hat{\mathcal{U}}_{L,\Delta}-\hat{\mathcal{A}}_{\Delta}$
where
$\hat{\mathcal{A}}_{\Delta}$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\Bigg{\{}I_{i}X_{i}^{T}\left(\hat{\beta}_{1,P}^{\Delta}-\hat{\beta}_{1,Q}^{\Delta}\right)-(1-I_{i})X_{i}^{T}\left(\hat{\beta}_{0,P}^{\Delta}-\hat{\beta}_{0,Q}^{\Delta}\right)\Bigg{\}},$
$\hat{\beta}_{j,P}^{\Delta}$ $=$ $\left(\sum_{i:I_{i}=j}X_{i}X_{i}^{T}\right)^{-1}\sum_{i:I_{i}=j}X_{i}\hat{P}^{\Delta}_{i},$
$\hat{\beta}_{j,Q}^{\Delta}$ $=$ $\left(\sum_{i:I_{i}=j}X_{i}X_{i}^{T}\right)^{-1}\sum_{i:I_{i}=j}X_{i}\hat{Q}^{\Delta}_{i}.$
Denote the adjusted $\Delta\hat{U}(\tau)$ by $\Delta\hat{U}^{(a)}(\tau)=\Delta\hat{U}(\tau)-\frac{n}{\sqrt{n_{0}}\sqrt{n_{1}}}\hat{\mathcal{A}}_{\Delta}$. Then $\hat{\mathcal{U}}_{CL,\Delta}=\frac{\sqrt{n_{0}}\sqrt{n_{1}}}{n}\{\Delta\hat{U}^{(a)}(\tau)-\Delta U(\tau)\}$. Define the covariate-adjusted test statistic on the difference of AUCs as $\mathcal{T}_{CL,\Delta}=\sqrt{n}\hat{\mathcal{U}}_{CL,\Delta}/\hat{\sigma}_{CL,\Delta}$, where $\hat{\sigma}_{CL,\Delta}^{2}=\hat{\sigma}_{L,\Delta}^{2}-\frac{n_{0}n_{1}}{n^{2}}(\hat{\beta}_{1,P}^{\Delta}-\hat{\beta}_{1,Q}^{\Delta}+\hat{\beta}_{0,P}^{\Delta}-\hat{\beta}_{0,Q}^{\Delta})^{T}\sum_{X}(\hat{\beta}_{1,P}^{\Delta}-\hat{\beta}_{1,Q}^{\Delta}+\hat{\beta}_{0,P}^{\Delta}-\hat{\beta}_{0,Q}^{\Delta})$. The theorem below establishes the asymptotic properties of $\hat{\mathcal{U}}_{CL,\Delta}$ and $\mathcal{T}_{CL,\Delta}$. The technical proofs are included in the Section B of the Supplementary Materials.
Theorem 1
Under the $H_{0,\Delta}:\Delta U(\tau)=0$ or the alternative hypothesis, $\sqrt{n}\hat{\mathcal{U}}_{CL,\Delta}\stackrel{{\scriptstyle d}}{{\to}}N\left(0,\sigma_{CL,\Delta}^{2}\right)$, $\hat{\sigma}_{CL,\Delta}^{2}\stackrel{{\scriptstyle p}}{{\to}}\sigma_{CL,\Delta}^{2}$, and $\mathcal{T}_{CL,\Delta}\stackrel{{\scriptstyle d}}{{\to}}N(0,1)$, where $\sigma_{CL,\Delta}^{2}=\sigma_{L,\Delta}^{2}-\frac{n_{0}n_{1}}{n^{2}}(\beta_{1,P}^{\Delta}-\beta_{1,Q}^{\Delta}+\beta_{0,P}^{\Delta}-\beta_{0,Q}^{\Delta})^{T}\sum_{X}(\beta_{1,P}^{\Delta}-\beta_{1,Q}^{\Delta}+\beta_{0,P}^{\Delta}-\beta_{0,Q}^{\Delta})$.
$H_{0,\Delta}$ is rejected when $|\mathcal{T}_{CL,\Delta}|>z_{\alpha/2}$ and the adjusted $100(1-\alpha)\%$ CI for $\Delta U(\tau)$ is $(\Delta\hat{U}^{(a)}(\tau)-z_{\alpha/2}SE\left\{\Delta\hat{U}^{(a)}(\tau)\right\},\Delta\hat{U}^{(a)}(\tau)+z_{\alpha/2}SE\left\{\Delta\hat{U}^{(a)}(\tau)\right\})$, where $SE\left\{\Delta\hat{U}^{(a)}(\tau)\right\}=\left(\frac{n}{n_{0}n_{1}}\hat{\sigma}_{CL,\Delta}^{2}\right)^{1/2}$.
As proved by Ye et al. [2024], $\sigma_{CL,\Delta}^{2}\leq\sigma_{L,\Delta}^{2}$, with strict inequality holding unless either (i) the covariates $X_{i}^{*}$ are uncorrelated with $P_{i}^{\Delta}$ and $Q_{i}^{\Delta}$, or (ii) the covariates not used in the randomization are uncorrelated with $P_{i}^{\Delta}$ and $Q_{i}^{\Delta}$, given the covariates used in the randomization. Therefore, the adjusted $\mathcal{T}_{CL,\Delta}$ provides a guaranteed efficiency gain over the unadjusted $\mathcal{T}_{L,\Delta}$.
2.2.2 Ratio of AUCs
Let $\hat{R}(\tau)=\hbox{log}\frac{\hat{U}_{1}(\tau)}{\hat{U}_{0}(\tau)}$ and $\hat{\mathcal{U}}_{L,R}=\frac{n_{0}n_{1}}{n}\left\{\hat{R}(\tau)-R(\tau)\right\}$. The linearization of $\hat{\mathcal{U}}_{L,R}$ is as follows:
@@ 2.2.2 Ratio of AUCs
where $\hat{P}^{R}_{i}=\left\{\frac{n_{0}}{\hat{U}_{1}(\tau)}\right\}^{I_{i}}\left\{\frac{n_{1}}{\hat{U}_{0}(\tau)}\right\}^{1-I_{i}}\hat{P}_{i}$ and $\hat{Q}^{R}_{i}=\left\{\frac{n_{0}}{\hat{U}_{1}(\tau)}\right\}^{I_{i}}\left\{\frac{n_{1}}{\hat{U}_{0}(\tau)}\right\}^{1-I_{i}}\hat{Q}_{i}$. The unadjusted test statistic on the ratio of AUCs is $\mathcal{T}_{L,R}=\sqrt{n}\hat{\mathcal{U}}_{L,R}/\hat{\sigma}_{L,R}$, where
@@
Derivations of $\hat{\mathcal{U}}_{L,R}$ and the asymptotic properties of $\mathcal{T}_{L,R}$ are provided in detail in the Section A.2 of the Supplementary Materials.
Similarly, we treat $\hat{P}^{R}_{i}$ and $\hat{Q}^{R}_{i}$ as outcomes and apply the generalized regression adjustment or augmentation [Tsiatis et al., 2008]. This results in the covariate-adjusted $\hat{\mathcal{U}}_{CL,R}$:
@@
Let $\hat{R}^{(a)}(\tau)$ denote the adjusted $\hat{R}(\tau)$, and is given by $\hat{R}^{(a)}(\tau)=\hat{R}(\tau)-\frac{n}{n_{0}n_{1}}\hat{\mathcal{A}_{R}}$. Then, $\hat{\mathcal{U}}_{CL,R}=\frac{n_{0}n_{1}}{n}\left\{\hat{R}^{(a)}(\tau)-R(\tau)\right\}$. Define the covariate-adjusted test statistic on the ratio of AUCs as $\mathcal{T}_{CL,R}=\sqrt{n}\hat{\mathcal{U}}_{CL,R}/\hat{\sigma}_{CL,R}$, where $\hat{\sigma}_{CL,R}^{2}=\hat{\sigma}_{L,R}^{2}-\frac{n_{0}n_{1}}{n^{2}}(\hat{\beta}_{1,P}^{R}-\hat{\beta}_{1,Q}^{R}+\hat{\beta}_{0,P}^{R}-\hat{\beta}_{0,Q}^{R})^{T}\sum_{X}(\hat{\beta}_{1,P}^{R}-\hat{\beta}_{1,Q}^{R}+\hat{\beta}_{0,P}^{R}-\hat{\beta}_{0,Q}^{R})$. The theorem below establishes the asymptotic properties of $\hat{\mathcal{U}}_{CL,R}$ and $\mathcal{T}_{CL,R}$, which can be proved similarly to the proof of Theorem 1.
@@
$H_{0,R}$ is rejected when $|\mathcal{T}_{CL,R}|>z_{\alpha/2}$ and the adjusted $100(1-\alpha)\%$ CI for $R(\tau)$ is $\big{(}\hbox{exp}\big{[}\hat{R}^{(a)}(\tau)-z_{\alpha/2}SE\left\{\hat{R}^{(a)}(\tau)\right\}\big{]},\hbox{exp}\big{[}\hat{R}^{(a)}(\tau)+z_{\alpha/2}SE\left\{\hat{R}^{(a)}(\tau)\right\}\big{]}\big{)}$, where $SE\left\{\hat{R}^{(a)}(\tau)\right\}=\left(\frac{n}{n_{0}^{2}n_{1}^{2}}\hat{\sigma}_{CL,R}^{2}\right)^{1/2}$. Similar to the discussion in Section 2.2.1, the adjusted $\mathcal{T}_{CL,R}$ provides a guaranteed efficiency gain over the unadjusted $\mathcal{T}_{L,R}$.

### v3

@@ Abstract
The area under the curve (AUC) of the mean cumulative function (MCF) has recently been introduced as a novel estimand for evaluating treatment effects in recurrent event settings, offering an alternative to the commonly used Lin-Wei-Yang-Ying (LWYY) model. The AUC of the MCF provides a clinically interpretable summary measure that captures the overall burden of disease progression, regardless of whether the proportionality assumption holds. To improve the precision of the AUC estimation while preserving its unconditional interpretability, we propose a nonparametric covariate adjustment approach. This approach guarantees efficiency gain compared to unadjusted analysis, as demonstrated by theoretical asymptotic distributions, and is universally applicable to various randomization schemes, including both simple and covariate-adaptive designs. Extensive simulations across different scenarios further support its advantage in increasing statistical power. Our findings highlight the importance of covariate adjustment for the analysis of AUC in recurrent event settings, offering practical guidance for its application in randomized clinical trials.
@@ 1 Introduction
In recent years, recurrent event analysis has gained traction in clinical studies, particularly in chronic conditions such as multiple hospitalizations for heart failure, recurrent relapses in multiple sclerosis, or frequent exacerbation in asthma [Solomon et al., 2024, Rogers et al., 2014, Cohen and Rudick, 2011, Keene et al., 2007]. A widely adopted approach for analyzing recurrent event data in clinical trials is the Lin-Wei-Yang-Ying (LWYY) model [Lin et al., 2000]. The LWYY model assumes proportionality between the mean cumulative functions (MCFs) of two arms, implying a constant rate ratio over time. This allows treatment effects to be evaluated through the ratio of the MCFs. In practice, this assumption might be violated, particularly in complex scenarios where treatment effects are time-varying. For example, the recent FINEARTS-HF study demonstrated time-dependent variation in rate ratio estimates [Vaduganathan et al., 2024]. While the LWYY model offers some robustness to model misspecification [Lin and Wei, 1989]—and its constant estimate can be interpreted as a weighted average of time-varying effects—this weighted average may lack clear clinical interpretability in some contexts, in which more flexible estimands that do not assume proportionality may be preferable.
The recent ICH E9(R1) addendum also emphasizes the importance of a robust definition quantifying the treatment effects in the settings of recurrent events with minimal assumptions. In this context, the area under the curve (AUC)-based approach for recurrent events proposed by Claggett et al. [2022] has gained considerable attention, offering a clinically meaningful summary of the cumulative burden of recurrent events without assuming proportionality [Gregson et al., 2023]. By calculating the area under the MCF for each treatment arm, one can derive a summary measure that reflects the overall mean burden of events experienced by patients over the follow-up period. A larger AUC signifies a greater cumulative burden, and treatment comparisons based on the difference or ratio of AUCs allow for intuitive interpretation in terms of absolute or relative reductions in disease burden. AUCs can be consistently estimated by the Ghosh-Lin estimator irrespective of whether the proportionality assumption holds, making it a robust and interpretable estimand for analyzing recurrent event data [Ghosh and Lin, 2000].
@@ 2 Area Under the Mean Cumulative Function
The area under the MCF was introduced [Claggett et al., 2022] to address the question: What is the cumulative burden of recurrent events over a fixed follow-up period? It quantifies this burden by capturing both the frequency and timing of recurrent events (e.g., repeated hospitalizations). Conceptually, each occurrence of a recurrent event increases the total time a subject spends in the undesirable condition, contributing to their cumulative burden, which is measured by event-time. The size of this contribution depends on timing: the earlier the event occurs, the more event-time it adds. The MCF tracks the average accumulation of such events across the study population. Therefore, the area under the MCF summarizes the mean total burden due to recurrent events, with a higher AUC indicating greater overall burden. This quantity provides a robust, model-free method for comparing treatment groups in terms of cumulative burden, with treatment effects expressed through differences or ratios of AUCs to reflect absolute or relative reductions in recurrent event burden. In what follows, we formally define the AUC estimand and describe how it can be estimated from censored recurrent event data.
@@ 2.1 Data and estimand
Let $D$ represent the time to a terminal event and write $N^{D}(t)=\mathbb{I}(D\leq t)$ as the counting process for $D$. Let $N^{*}(t)$ be the number of recurrent events up to time $t$. Let $C$ denote the independent censoring time. An observation consists of $\{N(\cdot),T,\delta\}$, where $N(t)=N^{*}(t\wedge C)$, $T=C\wedge D$, $\delta=\mathbb{I}(D\leq C)$, and $a\wedge b=\hbox{min}(a,b)$. The observed data set $\mathcal{D}=\{N_{i}(\cdot),T_{i},\delta_{i}\}$, for $i=1,\dots,n$, consists of independent realizations of $\{N(\cdot),T,\delta\}$.
Let $S(u)=\mathbb{P}(D\geq u)$ be the survival function, and $dR(u)=\mathbb{E}\{dN^{*}(u)\mid T\geq u\}=\mathbb{E}\{dN^{*}(u)\mid D\geq u\}$, where $dN^{*}(t)=\lim_{\Delta t\to 0}\{N^{*}(t+\Delta t^{-})-N^{*}(t^{-})\}$. Then, the mean cumulative function (MCF) is defined as
@@
The estimand $U(\tau)$ can be interpreted as the mean total event-free time lost from recurrent events over the interval $[0,\tau]$, which captures the cumulative burden of recurrent events in a clinically interpretable way. $U(\tau)$ is a natural extension of the restricted mean survival time (RMST) from time-to-first event settings to the recurrent event setting [Kim et al., 2017]. Specifically, in the time-to-first event setting, the MCF can be considered as the expected value of the counting process for the death event at time $t$, which is equivalent to the probability of experiencing death by time $t$—that is, 1 minus the survival probability at time $t$. As a result, in the time-to-first event setting, the area under the MCF over the interval $[0,\tau]$ equals $\tau$ minus the RMST over $[0,\tau]$.
Denote $U_{0}(\tau)$ and $U_{1}(\tau)$ as the area under the MCF for placebo and treatment arms. The treatment effect can be quantified as either the difference, $\Delta U(\tau)=U_{1}(\tau)-U_{0}(\tau)$, or the ratio, $U_{1}(\tau)/U_{0}(\tau)$. For convenience in later derivations, we define $R(\tau)=\hbox{log}\frac{U_{1}(\tau)}{U_{0}(\tau)}$, which represents the log-transformed ratio of the areas under the MCF. Similar to how the RMST is estimated based on the Kaplan-Meier estimator, the AUC under each treatment group can be nonparametrically estimated by plugging in the Ghosh-Lin estimator [Ghosh and Lin, 2000]. An open-source implementation of this method is available at https://github.com/zrmacc/MCC.
An estimand specifies five key attributes: the treatment (the specific intervention being evaluated), population (the group of subjects to which the treatment effect applies), variable (the outcome or endpoint being measured), intercurrent events (events occurring after randomization that may affect the interpretation or occurrence of the outcome), and the population-level summary (how the variable is summarized to compare treatment conditions). The five attributes of our estimand are defined in Table LABEL:tab:estimand.
2.2 Nonparametric Covariate Adjustment
We extend the nonparametric covariate adjustment approach introduced by Ye et al. [2022, 2023, 2024] to the AUC-based summary. The key concept involves linearizing the test statistic to derive an outcome for each patient, followed by the application of generalized regression adjustment or augmentation to these derived outcomes [Zhang et al., 2008, Tsiatis et al., 2008]. In this section, we focus on covariate adjustment for the ratio of AUCs. The corresponding method for the difference of AUCs follows analogously and is provided in Section A of the Supplementary Materials.
2.2.1 Preparation
Let $\hat{\mu}(t)$ and $\hat{U}(\tau)$ denote estimates of $\mu(t)$ and $U(\tau)$, obtained by substituting $S(u)$ with the Kaplan-Meier estimator $\hat{S}(u)$ and $dR(u)$ with the Nelson-Aalen estimator $d\hat{R}(u)=\sum_{i=1}^{n}dN_{i}(u)\big{/}\sum_{i=1}^{n}\mathbb{I}(T_{i}\geq u)$, where $N_{i}(t)=N_{i}^{*}(t\wedge C_{i})$ and $dN_{i}(t)=\lim_{\Delta t\to 0}\{N_{i}(t+\Delta t^{-})-N_{i}(t^{-})\}$.
@@ 2.2.1 Preparation
The estimates $\hat{\psi}_{i}(\tau)$, $\hat{P}_{i}$, and $\hat{Q}_{i}$ are obtained by replacing $S(u)$ with $\hat{S}(u)$, $\mu(s)$ with $\hat{\mu}(s)$, and estimating $\mathbb{P}(T\geq u)$ as $\sum_{i=1}^{n}\mathbb{I}(T_{i}\geq u)/n$. The martingales are estimated by: $d\hat{M}_{i}(u)=dN_{i}(u)-\mathbb{I}(T_{i}\geq u)d\hat{R}(u)$, and $d\hat{M}_{i}^{D}(u)=dN_{i}^{D}(u)-\mathbb{I}(T_{i}\geq u)d\hat{A}^{D}(u)$, with $d\hat{A}^{D}(u)=\sum_{i=1}^{n}dN^{D}_{i}(u)\big{/}\sum_{i=1}^{n}\mathbb{I}(T_{i}\geq u)$.
2.2.2 Covariate-Adjusted AUCs
@@
Let $\hat{R}(\tau)=\hbox{log}\frac{\hat{U}_{1}(\tau)}{\hat{U}_{0}(\tau)}$ and $\hat{\mathcal{U}}_{L,R}=\frac{n_{0}n_{1}}{n}\left\{\hat{R}(\tau)-R(\tau)\right\}$, where $n_{0}$ and $n_{1}$ represent the sample sizes for the placebo and treatment arms, respectively. The linearization of $\hat{\mathcal{U}}_{L,R}$ is as follows:
@@
where $I_{i}=1$ if subject $i$ is in the treatment arm and 0 otherwise, $\hat{P}^{R}_{i}=\left\{\frac{n_{0}}{\hat{U}_{1}(\tau)}\right\}^{I_{i}}\left\{\frac{n_{1}}{\hat{U}_{0}(\tau)}\right\}^{1-I_{i}}\hat{P}_{i}$ and $\hat{Q}^{R}_{i}=\left\{\frac{n_{0}}{\hat{U}_{1}(\tau)}\right\}^{I_{i}}\left\{\frac{n_{1}}{\hat{U}_{0}(\tau)}\right\}^{1-I_{i}}\hat{Q}_{i}$. The unadjusted test statistic on the ratio of AUCs is $\mathcal{T}_{L,R}=\sqrt{n}\hat{\mathcal{U}}_{L,R}/\hat{\sigma}_{L,R}$, where
@@
Here, $\hat{\psi}_{1}(\tau)$ and $\hat{\psi}_{0}(\tau)$ consist of $\hat{\psi}_{i}(\tau)$ from subjects in the treatment and placebo arms, respectively. Derivations of $\hat{\mathcal{U}}_{L,R}$ and the asymptotic properties of $\mathcal{T}_{L,R}$ are provided in detail in Section B of the Supplementary Materials.
Let $\boldsymbol{X}^{*}$ be a $p$-dimensional vector of baseline covariates. Define $X_{i}=X_{i}^{*}-\bar{X}$, where $X_{i}^{*}$ is the $i$th column of $\boldsymbol{X}^{*}$, and $\bar{X}$ is the sample mean of all the $X_{i}^{*}$, ensuring $\mathbb{E}(X_{i})=0$. Following Ye et al. [2024], we treat $\hat{P}^{R}_{i}$ and $\hat{Q}^{R}_{i}$ as outcomes and apply the generalized regression adjustment or augmentation [Tsiatis et al., 2008]. This results in the covariate-adjusted $\hat{\mathcal{U}}_{CL,R}$:
@@
Let $\hat{R}^{(a)}(\tau)$ denote the adjusted $\hat{R}(\tau)$, and is given by $\hat{R}^{(a)}(\tau)=\hat{R}(\tau)-\frac{n}{n_{0}n_{1}}\hat{\mathcal{A}_{R}}$. Then, $\hat{\mathcal{U}}_{CL,R}=\frac{n_{0}n_{1}}{n}\left\{\hat{R}^{(a)}(\tau)-R(\tau)\right\}$. Define the covariate-adjusted test statistic on the ratio of AUCs as $\mathcal{T}_{CL,R}=\sqrt{n}\hat{\mathcal{U}}_{CL,R}/\hat{\sigma}_{CL,R}$, where $\hat{\sigma}_{CL,R}^{2}=\hat{\sigma}_{L,R}^{2}-\frac{n_{0}n_{1}}{n^{2}}(\hat{\beta}_{1,P}^{R}-\hat{\beta}_{1,Q}^{R}+\hat{\beta}_{0,P}^{R}-\hat{\beta}_{0,Q}^{R})^{T}\sum_{X}(\hat{\beta}_{1,P}^{R}-\hat{\beta}_{1,Q}^{R}+\hat{\beta}_{0,P}^{R}-\hat{\beta}_{0,Q}^{R})$. The theorem below establishes the asymptotic properties of $\hat{\mathcal{U}}_{CL,R}$ and $\mathcal{T}_{CL,R}$. The technical proofs are included in Section C of the Supplementary Materials.
@@
$H_{0,R}$ is rejected when $|\mathcal{T}_{CL,R}|>z_{\alpha/2}$ and the adjusted $100(1-\alpha)\%$ CI for $R(\tau)$ is $\big{(}\hbox{exp}\big{[}\hat{R}^{(a)}(\tau)-z_{\alpha/2}SE\left\{\hat{R}^{(a)}(\tau)\right\}\big{]},\hbox{exp}\big{[}\hat{R}^{(a)}(\tau)+z_{\alpha/2}SE\left\{\hat{R}^{(a)}(\tau)\right\}\big{]}\big{)}$, where $SE\left\{\hat{R}^{(a)}(\tau)\right\}=\left(\frac{n}{n_{0}^{2}n_{1}^{2}}\hat{\sigma}_{CL,R}^{2}\right)^{1/2}$.
As proved by Ye et al. [2024], $\sigma_{CL,R}^{2}\leq\sigma_{L,R}^{2}$, with strict inequality holding unless either (i) the covariates $X_{i}^{*}$ are uncorrelated with $P_{i}^{R}$ and $Q_{i}^{R}$, or (ii) the covariates not used in the randomization are uncorrelated with $P_{i}^{R}$ and $Q_{i}^{R}$, given the covariates used in the randomization. Therefore, the adjusted $\mathcal{T}_{CL,R}$ provides a guaranteed efficiency gain over the unadjusted $\mathcal{T}_{L,R}$.